Circular No. 7172

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) BMARSDEN@CFA.HARVARD.EDU or DGREEN@CFA.HARVARD.EDU (science)

        URL http://cfa-www.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7244/7440/7444 (for emergency use only)

SUPERNOVA 1999cd IN NGC 3646 P. Garnavich, S. Jha, R. Kirshner, and P. Challis, Harvard-Smithson-

ian Center for Astrophysics, report that a spectrum of SN 1999cd was obtained by P. Berlind with the 1.5-m Tillinghast telescope on May 15.3 UT. The spectrum shows a broad Hα feature with a P-Cyg profile, indicating that the object is a type-II supernova. Weaker He I 587.5-nm and Hβ features are also present. The narrow emission lines from the host galaxy give a recession velocity of 4260 km/s. The photospheric velocity of the supernova, derived from the minimum of the Hα absorption, is 14 000 km/s, but the absorption extends to nearly 20 000 km/s.

F. Patat, European Southern Observatory (ESO); L. Rizzi, Universitá di Padova; and L. Guzzo and G. Guerrero, Osservatorio Astronomico di Brera, communicate: “We have observed SN 1999cd with the ESO 3.6-m telescope (+ EFOSC2) on May 19.09 UT at La Silla. A full reduction of the CCD spectrogram (range 340–740 nm, resolution 1.8 nm) shows spectral features (very-well-developed P-Cyg profiles and a rather red continuum) that are consistent with a supernova of type II, a few weeks after maximum light. The most prominent features are identified as Hα, Hβ, Hγ, Na I D, Fe I, Ba II, Ti II, and Ca II. The expansion velocities deduced from the minima of Hα and Ca II H and K are 13 600 and 9200 km/s, respectively. Hα shows a boxy absorption profile and a rather flat-topped emission, with some possible structures. The recession velocity of NGC 3646 at the supernova’s location, deduced from the narrow Hα probably coming from an underlying H II region, is 4390 km/s. The magnitude of SN 1999cd is R ∼ 17.4.”

GRB 990510 S. Covino, Osservatorio Astronomico di Brera, reports on behalf of

a large collaboration (including D. Lazzati, G. Ghisellini, P. Saracco, S. Campana, G. Chincarini, S. Di Serego, A. Cimatti, L. Vanzi, L. Pasquini, F. Haardt, M. Vietri, L. Stella, H. Boehnhardt, F. Bresolin, P. Moller, and G. Rupprecht): “We made R-band imaging polarimetry of the optical transient associated to GRB 990510 at the Very Large Telescope of the European Southern Observatory on May 11.13509 UT, when object was at R = 19.1 ± 0.02. We found linear polarization at the level of 1.7 ± 0.2 percent with p.a. 11^{o} ± 3^{o}, relative to the stars in the field.”

1999 May 19  \begin{matrix}c \\ ◯ \end{matrix} Copyright 1999 CBAT  Daniel W. E. Green